# Zero-shot Class Unlearning via Layer-wise Relevance Analysis and Neuronal Path Perturbation

### v1

@@ V-D2 Different Amount of Perturbed Neurons
After analyzing the same number of neurons, perturbing different quantities of neurons can also impact unlearning performance.
Table 6(a) demonstrates the performance of our unlearning method on the MNIST dataset when analyzing the top 50 neurons’ relevance. Specifically, for $r=\text{epsilon}$, as the unlearning percentage $M_{p}$ increases, the accuracy $A_{t}$ decreases steadily. At $M_{p}=16\%$, the value of $A_{t}$ drops significantly to $0.005$, indicating that the unlearning process stabilizes and effectively erases the targeted information. Beyond this point, further increases in $M_{p}$ do not result in a significant change, as seen with $M_{p}=20\%$, where $A_{t}$ reaches $0.001$.
On the CIFAR-10 dataset, using the epsilon rule, $A_{t}$ (target class accuracy) shows a significant decrease as $M_{p}$ increases, while $A_{g}$ (non-target class accuracy) remains stable around $0.98$ or $0.97$. In Table 6(c), specifically, at $M_{p}=0\%$, $A_{t}=0.99$, indicating near-perfect accuracy for the target class. As $M_{p}$ increases to $10\%$ and $15\%$, $A_{t}$ drops sharply to $0.67$ and $0.34$. At $M_{p}=20\%$, $A_{t}$ further declines to $0.11$, and finally reaches its lowest values of $0.03$ and $0.06$ at $M_{p}=25\%$ and $30\%$, respectively. This trend demonstrates that as $M_{p}$ increases, the target class accuracy decreases progressively, achieving a strong unlearning effect.
Table 6(e) demonstrates the performance of applying our method to the CIFAR-100 dataset. Under the epsilon rule, $A_{t}$ decreases gradually as $M_{p}$ increases, while $A_{g}$ remains constant at $0.83$. Specifically, when $M_{p}$ increases from $0\%$ to $2\%$, $A_{t}$ slightly decreases from $0.88$ to $0.85$. As $M_{p}$ reaches $4\%$, $A_{t}$ drops more significantly to $0.56$, and further decreases to $0.18$ at $M_{p}=8\%$. Finally, at $M_{p}$ values of $10\%$ and $12\%$, $A_{t}$ reaches $0.00$, achieving complete unlearning.
A similar trend is observed in Table 6(h), which analyzes the top 150 neurons’ relevance on the mini-ImageNet dataset. For $r=\text{epsilon}$, the accuracy $A_{t}$ declines steadily as $M_{p}$ increases. When $M_{p}$ reaches $10\%$, $A_{t}$ drops to $0.00$, and even with further increases in $M_{p}$, this minimal accuracy remains, indicating that unlearning has effectively stabilized.
The trend is consistent across all relevance propagation rules $r$: as $M_{p}$ increases, $A_{t}$ decreases significantly, reaching a point where further unlearning does not substantially affect the results, indicating prosperous and stable unlearning.
Perturbing only a small subset of neurons fails to significantly affect a model’s memory because of the model’s redundancy. Information is often encoded across multiple neurons and layers in neural networks, especially those trained on complex tasks. This redundancy means that specific data points or features are not isolated to just a few neurons; their memory is distributed throughout the network.
Our experimental results reinforce this idea. For example, when analyzing the top $50$ neurons on the MNIST dataset or the top $150$ neurons on the mini-ImageNet dataset, significant memory disruption only occurs when a sufficient percentage of neurons are affected ($M_{p}\geq 16\%$ for MNIST and $M_{p}\geq 10\%$ for mini-ImageNet). This indicates that perturbing a small portion of neurons is insufficient to overcome the model’s redundancy, as the remaining unaffected neurons can compensate, preserving the model’s memory.
@@ V-D3 Different Amount of Analyzed Neurons
Analyzing different quantities of neurons related to unlearning data can impact unlearning performance.
When comparing Table 6(a) and Table 6(b), we can observe that when analyzing the top $50$ neurons, unlearning is effectively achieved at $M_{p}$ = $16\%$ for most methods, with a further increase in $M_{p}$ to $20\%$ improving the unlearning effect. However, when analyzing the top 100 neurons, similar unlearning results are only reached when $M_{p}$ is $20\%$ or $24\%$.
Specifically, when analyzing the top $50$ neurons, at $M_{p}$ = $16\%$, utilizing propagation rules like epsilon, gamma+epsilon, alpha1beta0, alpha2beta1, and PatternAttribution $S(x){+-}$ achieve successful unlearning, with the target accuracy ($A_{t}$) dropping close to 00, while the global accuracy ($A_{g}$) remains high. In contrast, when analyzing the top $100$ neurons, only the PatternAttribution $S(x){+-}$ method achieves effective unlearning at $M_{p}$ = $16\%$, while other methods require $M_{p}$ to reach $20\%$ or more to achieve a similar drop in $A_{t}$, even though $A_{g}$ stays consistent.
When the number of neurons analyzed increases to $100$, while it includes more neurons related to the unlearning data, it also introduces more weak or unrelated neurons. These unrelated neurons can interfere with the method’s statistical steps, making it harder to identify the neurons relevant to the unlearning data accurately. Therefore, when we analyze more neurons, a higher $M_{p}$ is needed to achieve effective unlearning.
We also compare Table 6(c) and Table 6(d), which demonstrate the performance of unlearning on the CIFAR-10 dataset by analyzing the top 150 and 200 relevance. In Table 6(c), which analyzes the relevance of the top 150 neurons, we observe that the unlearning effect becomes prominent as the $M_{p}$ increases. At $M_{p}=25\%$, the $A_{t}$ drops sharply to $0.03$, and further to $0.06$ at $M_{p}=30\%$, while the $A_{g}$ remains stable at $0.97$. These bolded values indicate that the unlearning requirements are effectively met, as the target test accuracy is significantly reduced while the generalization accuracy stays consistent. This shows a clear unlearning effect when removing relevant neurons from the decision process at higher dropout probabilities.
In Table 6(d), where the top 200 neurons are analyzed, the unlearning effect becomes even more pronounced. At $M_{p}=25\%$, $A_{t}$ drops to $0.01$, and at $M_{p}=30\%$, the accuracy on the target test set reaches $0.00$, signifying near-complete unlearning. Meanwhile, the $A_{g}$ decreases slightly to $0.95$. Comparing the two tables, the unlearning effect intensifies as the number of neurons being analyzed increases. This suggests that analyzing more neurons allows for more thorough removal of target class information, leading to a stronger unlearning outcome.
In Table 6(e), which analyzes the relevance of the top 150 neurons on the CIFAR-100 dataset, the unlearning effect becomes evident as the $M_{p}$ increases. When $M_{p}=10\%$, the $A_{t}$ drops sharply to $0.01$, while the $A_{g}$ remains stable at $0.81$. This shows that the method achieves the desired unlearning at this point. At $M_{p}=12.5\%$, the target test accuracy reaches $0.00$, indicating complete unlearning of the target class, while $A_{g}$ still holds at $0.81$. These bolded values highlight that unlearning requirements are successfully met, as the target test accuracy significantly decreases with minimal impact on global accuracy.
Similarly, the results are consistent in Table 6(f), which considers the top 200 neurons. At $M_{p}=10\%$, $A_{t}$ drops to $0.01$, and at $M_{p}=12.5\%$, it reaches $0.00$, signifying a full unlearning effect. The generalization accuracy ($A_{g}$) remains at $0.81$ throughout, indicating that unlearning is effectively isolated to the target class without disrupting the overall performance. Comparing the two tables, increasing the number of neurons analyzed leads to a consistent unlearning trend, with unlearning performance stabilizing as more neurons are involved.
When comparing Table 6(g), Table 6(h), and Table 6(i), which examine the performance of unlearning on the mini-ImageNet dataset by analyzing the top 100, 150, and 200 relevance, we can observe some significant differences. The experimental results using other rules for $r$ show similar trends to those observed when $r$ is set to epsilon. Therefore, we will mainly focus on the experimental results and their changes when using epsilon as a propagation rule.
In Table 6(g), where the top $100$ neurons’ relevance is analyzed using $r=$ epsilon, the method begins to show an unlearning effect at $M_{p}=4\%$, with a noticeable drop in target accuracy $A_{t}$ to $0.61$. As $M_{p}$ increases to $8\%$, $A_{t}$ further decreases to $0.32$, and at $M_{p}=10\%$, it drops significantly to $0.08$. By $M_{p}=12\%$, $A_{t}$ stabilizes at $0.08$, while the global accuracy $A_{g}$ remains at $0.83$.
Building on this, in Table 6(h), the unlearning effect is more pronounced even at a lower $M_{p}$ when we analyze the top $150$ neurons with $r=$ epsilon. Here, $A_{t}$ drops to $0.56$ at $M_{p}=4\%$. As $M_{p}$ reaches $8\%$, $A_{t}$ decreases to $0.18$. At both $M_{p}=10\%$ and $M_{p}=12\%$, $A_{t}$ is reduced to 00, indicating complete unlearning, while $A_{g}$ remains consistent at $0.83$.
Similarly, in Table 6(i), where the top $200$ neurons’ relevance is analyzed with $r=$ epsilon, the trends are similar to those observed with the top $150$ neurons. Unlearning starts at $M_{p}=4\%$ with $A_{t}$ dropping to $0.56$, and as $M_{p}$ increases to $8\%$, $A_{t}$ falls to $0.18$. At both $M_{p}=10\%$ and $M_{p}=12\%$, $A_{t}$ reaches 00, while $A_{g}$ remains stable at $0.83$.
These results indicate that increasing the number of neurons analyzed from $100$ to $150$ leads to more effective unlearning at slightly lower $M_{p}$ values. However, increasing the number of neurons to $200$ does not enhance the unlearning effect beyond what is achieved with $150$ neurons. This pattern holds true across all relevance rules, demonstrating that the specific choice of $r$ has minimal impact on the method’s overall effectiveness.
@@ V-E1 Impact of Neural Network Redundancy on the Number of Perturbed Neurons
Model parameter redundancy refers to the presence of a large number of parameters in deep learning models, where some parameters may not be essential for the model’s performance. This redundancy is especially notable in deep learning models, which are typically over-parameterized, meaning the number of parameters far exceeds the amount of training data needed. The redundancy also presents certain challenges when implementing machine unlearning.
When considering models trained on datasets with different numbers of classes but identical parameter counts, an interesting observation arises. Models trained on datasets with fewer classes, such as MNIST or CIFAR-10, tend to have higher redundancy. In these cases, it may require perturbing $16\%$ to $20\%$ of neurons in the fully connected layers to achieve the goal of unlearning.
Many parameters are redundant in models with high redundancy and contribute less to the model’s performance. Therefore, to forget specific information, it may be necessary to perturb a larger number of neurons to remove all relevant information. In other words, high redundancy increases the number of neurons that need to be perturbed, as individual perturbations may not significantly affect the model’s overall performance.
On the other hand, models trained on datasets with more classes, like CIFAR-100 or mini-ImageNet, exhibit lower redundancy. In such models, perturbing only $10\%$ of neurons in the fully connected layers is often sufficient for unlearning. This happens because each parameter becomes more critical to performance with more classes, making the model more sensitive to even small perturbations.
Each parameter contributes more significantly to the model’s performance in models with low redundancy. Since each parameter and neuron plays a more crucial role in such models, perturbing these parameters will directly impact the model’s behavior and memory. Consequently, in low redundancy models, perturbing a small number of neurons can substantially alter the model’s memory and achieve the goal of machine unlearning. This implies that low redundancy models may be more efficient in implementing machine unlearning because fewer perturbations are needed.
@@ V-E2 Impact of Neural Network Redundancy on the Model Utility
Unlearning methods, particularly those targeting the perturbation of specific neuron weights within a neural network, have been shown to have minimal adverse effects on overall model utility. This phenomenon can be attributed to the inherent redundancy present within neural networks. Neural networks are designed with multiple layers and numerous neurons per layer, enabling them to distribute the task of data inference across a vast array of pathways. Consequently, the interference with a small subset of neuron weights does not singularly disrupt the network’s functionality.
The redundancy in neural networks means that no single pathway is solely responsible for the model’s ability to infer and classify data. Instead, the model’s predictive performance arises from the collective interactions of numerous neurons and pathways. This architectural characteristic ensures that the perturbation of a few specific neurons or parameters does not drastically affect the model’s capacity to generalize from the data. Such robustness is crucial, especially when considering the need for selective unlearning or model editing techniques to remove specific information or biases.
Additionally, neurons within a neural network exhibit differential sensitivity to various data types. The neurons activated by one input type may differ from those triggered by another. This differential activation pattern implies that altering the weights associated with specific neurons will not uniformly impact the model’s utility across all data types. For instance, modifying weights connected to neurons predominantly involved in recognizing features specific to one class of data may leave the recognition capabilities for other classes largely unaffected.
@@ VI Conclusion
Experimental results on image classification models demonstrate the effectiveness of the proposed method in removing targeted data while maintaining overall model performance. Our method contributes to privacy-preserving machine learning and provides a practical solution for handling data deletion requests and safeguarding user privacy. Future work may focus on further optimizing the method and exploring its applications in more complex scenarios.

### v2

@@ V-D2 Different Amount of Perturbed Neurons
Our experiments, employing the epsilon rule within the layer-wise relevance propagation framework, investigated the effect of varying the proportion of perturbed neurons ($M_{p}$) on unlearning performance across different datasets.
Consistent across MNIST (Table VII), CIFAR-10 (Table VIII), CIFAR-100 (Table IX), and mini-ImageNet (Table X), we observed a general trend: as $M_{p}$ increases, the target accuracy ($A_{t}$) steadily declines. Notably, $A_{t}$ eventually reaches a stable low point, beyond which further perturbation yields minimal additional unlearning. For instance, on MNIST, $A_{t}$ stabilizes around $0.005$ at $M_{p}=16\%$, with only marginal reduction at $20\%$. Similar stabilization trends are evident on CIFAR-10 (around $M_{p}=25\%-30\%$), CIFAR-100 ($M_{p}\geq 10\%$), and mini-ImageNet ($M_{p}\geq 10\%$). Throughout this process, the non-target accuracy ($A_{g}$) remains relatively high, indicating preserved model utility.
This behavior suggests that perturbing a small fraction of neurons does not drastically impair the model’s overall memory capabilities. We attribute this resilience to the inherent redundancy in neural network representations, where information is distributed across numerous neurons and layers. Consequently, perturbing a limited number of neurons is often insufficient to overcome this redundancy, as the unperturbed neurons can compensate for the induced disruptions. Our findings support this hypothesis, as significant memory disruption, indicated by substantial and stable drops in $A_{t}$, typically occurs only when $M_{p}$ reaches a certain threshold (e.g., $\geq 16\%$ for MNIST, $\geq 10\%$ for mini-ImageNet). This demonstrates that a critical mass of perturbed neurons is required to effectively break through the distributed nature of learned information and achieve stable unlearning.
@@ V-D3 Different Amount of Analyzed Neurons
The analysis of different quantities of neurons identified as relevant to the unlearning target significantly impacts the effectiveness of the unlearning process, as demonstrated by the trends in $A_{g}$ and $A_{t}$ across varying $M_{p}$ and $N$.
Examining the results on MNIST (Table VII), we observe a clear influence of the number of analyzed neurons, top $N$. For the first four rules (epsilon, gamma+epsilon, alpha1beta0, alpha2beta1), when analyzing the top 50 neurons, a substantial drop in $A_{t}$ occurs rapidly as $M_{p}$ increases beyond 8%, reaching near zero values by $M_{p}=16\%$. However, when analyzing the top 100 neurons, a comparable level of $A_{t}$ reduction is only achieved at higher perturbation levels, typically at $M_{p}=20\%$. This suggests that for MNIST, increasing the number of analyzed neurons beyond a certain point ($N=50$ in this case) might include less relevant neurons. As a result, a higher $M_{p}$ is needed across the entire larger group of neurons to achieve the desired unlearning effect, compared to focusing on a smaller, more critical subset of neurons.
The relationship between $N$ and unlearning effectiveness appears dataset-dependent. On CIFAR-10 (Table VIII) and mini-ImageNet (Table X), increasing $N$ seems beneficial up to a certain point. For CIFAR-10, comparing $N=150$ and $N=200$ under the four rules, $N=200$ leads to a more complete reduction in $A_{t}$ at the highest $M_{p}$ values (e.g., $A_{t}$ reaches 0.00 at $M_{p}=30\%$ for $N=200$, whereas it’s around 0.06 for $N=150$). Similarly, on mini-ImageNet, increasing $N$ from 100 to 150 leads to a faster and more complete $A_{t}$ drop, reaching 0.00 by $M_{p}=10\%$, compared to 0.08 for $N=100$. However, further increasing $N$ from 150 to 200 on mini-ImageNet does not yield significant additional improvement in $A_{t}$ reduction at high $M_{p}$. These findings suggest there may be an optimal number of neurons to analyze for unlearning, where increasing $N$ initially helps capture more relevant information, but exceeding this optimum introduces noise or targets neurons less crucial for the unlearning task.
On CIFAR-100 (Table IX), within the tested range of $N$ (150 and 200), both values exhibit very similar unlearning characteristics under all rules. The $A_{t}$ values drop sharply to nearly zero by $M_{p}=10\%$ for both $N=150$ and $N=200$, with $A_{g}$ remaining stable. For CIFAR-100, within this range, the number of analyzed neurons has less impact on the unlearning performance compared to the level of $M_{p}$.
When selecting the number of neurons to analyze, there exists a critical balance point. Analyzing too few neurons may fail to capture the truly strongly relevant neurons, resulting in incomplete unlearning. Conversely, analyzing too many neurons can introduce many irrelevant or weakly related neurons, creating noise and reducing the efficiency of the unlearning process. Therefore, finding an appropriate value of $N$—balancing the capture of key neurons and the avoidance of noise—is essential for achieving effective and efficient unlearning.
@@ V-E1 Impact of Neural Network Redundancy on the Number of Perturbed Neurons
Deep learning models, particularly over-parameterized ones, often exhibit significant parameter redundancy. This means the model contains more parameters than strictly necessary to achieve a given level of performance on the training data, leading to information being distributed across many neurons and pathways. This inherent redundancy can influence how effectively specific information can be unlearned by perturbing a subset of neurons.
Our experimental results suggest that the degree of unlearning required, measured by $M_{p}$, is related to the apparent redundancy in the model’s representation of the target class, which may vary with dataset complexity. For datasets with fewer classes, such as MNIST and CIFAR-10, where models might develop more redundant representations for each class, achieving effective unlearning typically required perturbing a higher proportion of neurons, with $A_{t}$ dropping significantly only at $M_{p}$ values ranging from 16% to 20% (for MNIST) or higher (for CIFAR-10).
This necessity for higher $M_{p}$ in seemingly more redundant representations can be attributed to the distributed nature of the learned information. When knowledge about a target class is spread across a larger number of parameters or neurons, perturbing only a small fraction of them has a limited impact on the overall representation of that class. Therefore, a larger proportion of neurons must be perturbed to effectively disrupt the redundant encoding and significantly reduce the model’s ability to recognize the target class.
In contrast, for datasets with a larger number of classes, such as CIFAR-100 and mini-ImageNet, where the model might develop less redundant or more specialized representations due to the complexity of the task, effective unlearning was achieved at lower $M_{p}$ values. In these cases, perturbing around 10% of neurons was often sufficient to cause a drastic drop in $A_{t}$. This observation supports the idea that in less redundant representations, the information might be more concentrated, making the model more sensitive to perturbations of fewer, more critical neurons.
@@ V-E2 Impact of Neural Network Redundancy on the Model Utility
TABLE XI: Analysis of the Overlap of Relevant Neurons in the Perturbed Layer
$D$ Top $N$ $M_{p}~{}(\approx)$ $\mathcal{X}_{most}$@2 $\mathcal{X}$@10 $Share$@$most$
Proportion Quantity Proportion Quantity Proportion Quantity Quantity Position
MNIST 50 0.16 80 0.43 34 0 0 6 [155, 345]
CIFAR-10 200 0.20 400 0.45 178 0 0 7 [1447]
CIFAR-100 150 0.10 200 0.30 60 0 0 5 [1064]
mini-ImageNet 150 0.10 400 0.47 189 0 0 9 [3452]
A key finding from our experiments is that the neuron perturbation unlearning methods generally have minimal impact on the model’s overall utility, as measured by $A_{g}$. This robustness of $A_{g}$ during targeted unlearning is largely a consequence of the inherent redundancy present in deep neural networks. In these over-parameterized models, data inference and feature extraction are distributed across numerous layers, neurons, and intricate pathways [26, 27].
We conducted experiments on the layers of different neural networks trained on four datasets to analyze the distribution of class-related information. We selected the top ten classes from four datasets and conducted experimental analysis under the condition of minimizing the parameters required for unlearning. We defined the following metrics to analyze the information overlap between neurons:
  • $\mathcal{X}_{most}$@2 represents the maximum number of shared neurons among strongly relevant neurons in the perturbed FC layer between any two classes. The more shared relevant neurons required to store information for both classes, the more likely unlearning one class will affect the model’s classification utility for the other class.
  • $\mathcal{X}$@10 represents the number of shared neurons in the FC layer perturbed by our method that exhibit strong relevance with the classification of ten classes. A higher value indicates that path perturbation on one class is more likely to adversely affect the model’s classification utility for other classes.
  • $Share$@$most$ indicates the maximum number of classes whose classifications are correlated with a single neuron among the strongly correlated neurons we analyzed.
In Table XI, we observe that the overlap of neurons strongly relevant to two different classes in a single FC layer of neural networks trained on four datasets, $\mathcal{X}_{most}$@2, is less than $0.5$. This phenomenon indicates that, although neurons in a single layer do share information for certain classes, each class has its own strongly correlated neurons, allowing perturbation of one class’s classification path without affecting the model’s utility for other classes.
Furthermore, we aimed to verify whether any neurons are strongly relevant to all ten classes. As shown in Table XI, the experimental results indicate that $\mathcal{X}$@10 is 00 for all models upon completing unlearning, meaning no single neuron simultaneously serves the classification of these ten classes. This further confirms the specific relevance of neurons with different class data.
Does a single fully connected layer (FC layer) in a neural network contain neurons strongly relevant to multiple classes? We further analyzed this by calculating the $Share$@$most$ for each model. We observed that, when analyzing the strongly relevant neurons for the ten classes in the MNIST dataset, only two neurons (No. $135$ and No. $345$) were shared by a maximum of six classes. In two ResNet-50 models trained on CIFAR-10 and CIFAR-100, only one neuron each (No. $1447$ and No. $1064$) was shared by a maximum of seven and five classes, respectively. This result suggests that the knowledge learned for each class is represented by specific strongly relevant neurons forming distinct classification paths. For VGG-16 trained on the mini-ImageNet dataset, which has a single layer with $4096$ neurons, we calculated the top 400 strongly relevant neurons for each of the ten classes, but found only one neuron (No. 3452) serving a maximum of nine classes’ classification tasks.
The low neuron overlap, absence of universally critical neurons demonstrate that neural networks have sufficient redundancy to maintain distinct classification paths for each class. Due to this redundancy, no single neuron or a small set of neurons is solely responsible for the model’s entire performance or the inference for general classes. Instead, the model’s ability to generalize relies on the collective computation and interaction of a vast number of parameters. Consequently, perturbing or modifying a limited subset of neurons specifically identified for unlearning the target class does not cause a catastrophic breakdown of the network’s fundamental functionality or its ability to correctly classify examples from other classes.